the number of five digits numbers that can be formed without any restriction is
a. 990000
b. 100000
c. 90000
d. none of these
please need the procedure

Grade:12

6 Answers

jitender lakhanpal
10 years ago

hi shuvam

option c

 

without any restriction means without repitition

9 * 10 * 10 * 10 * 10  = 90000

use fundamental theorom of addition and multiplication

in ten thousands place you can place 9 digit ( excluding 0 ) in all the other places we can place 10 digits

aa aa
10 years ago

Each digit can be 0 or 1 or 2 .. or 10.

10 choices * 10 choices * ... 10 choices = 100000

HTH

srinivas aytha
10 years ago
9ways 10ways 10ways 10ways 10ways

9xx10x10x10x10=90000 five digit numbers with out any restriction

sridhar yelkur
10 years ago

ANSWER IS OBVIOUSLY,,,,,,,,9*10*10*10*10=90000

GOPI
10 years ago
If u say each distinct digit. There are total 10 digits including "0".I need to prepare 5 digits number.so I have 10p5 permutations with out repetition.but in them there are such numbers starting with 0.so total numbers of that kind 9p4.so ans is 10p5-9p4. ..... else ..... 9*10*10*10*10 ways obviously ..........................................
venkatesh vadde
10 years ago
They are right the answer is 9*10*10*10*10=90000 option (C)